# Training and Evaluating Language Models with Template-based Data Generation

### v1

@@ Abstract
The rapid advancement of large language models (LLMs) such as GPT-3, PaLM, and Llama has significantly transformed natural language processing, showcasing remarkable capabilities in understanding and generating language. However, these models often struggle with tasks requiring complex reasoning, particularly in mathematical problem-solving, due in part to the scarcity of large-scale, high-quality, domain-specific datasets necessary for training sophisticated reasoning abilities. To address this limitation, we introduce Template-based Data Generation (TDG), a novel approach that leverages LLMs (GPT-4) to automatically generate parameterized meta-templates, which are then used to synthesize a vast array of high-quality problems and solutions. Leveraging TDG, we create TemplateMath Part I: TemplateGSM, a dataset comprising over 7 million synthetically generated grade school math problems—each accompanied by code-based and natural language solutions—with the potential to generate an effectively unlimited number more. This dataset alleviates the scarcity of large-scale mathematical datasets and serves as a valuable resource for pre-training, fine-tuning, and evaluating LLMs in mathematical reasoning. Our method not only enables the generation of virtually infinite data but also elevates data augmentation to a new level by using GPT-4 for meta-template generation, ensuring diverse and high-quality problem structures. The TemplateMath Part I: TemplateGSM dataset is publicly available at https://huggingface.co/datasets/math-ai/TemplateGSM.111The code is available at https://github.com/iiis-ai/TemplateMath.
@@ 1 Introduction
We introduce TDG, a scalable method for generating an effectively infinite amount of high-quality, domain-specific data using parameterized templates generated by GPT-4.
By leveraging GPT-4 to create meta-templates, we advance data augmentation to a new level, ensuring a diverse and rich set of problem structures for data synthesis.
@@ 2 Template-based Data Generation
Template-based Data Generation (TDG) is a method designed to systematically produce a vast array of mathematical problems along with their corresponding solutions by leveraging parameterized templates. To elevate data augmentation to a new level, we employ GPT-4 to generate these meta-templates, capturing a wide variety of problem structures and linguistic styles. By varying parameters within these GPT-4-generated templates, TDG ensures both scalability and quality in the generated data. This approach enables the creation of diverse and complex problem sets, which are essential for training and evaluating large language models in mathematical reasoning tasks.
@@ 2.1 Methodology
The TDG process involves several key components that work together to generate high-quality mathematical datasets:
@@ 2.1.1 Generation of Meta-Templates with LLMs
We begin by utilizing large language models (LLMs), such as GPT-4, to generate meta-templates (Zhang et al., 2023) that capture the underlying structures of various mathematical problem types. GPT-4’s advanced language generation capabilities allow us to produce a diverse set of templates encompassing a wide range of mathematical concepts and problem types. These templates include placeholders for variable components such as names, items, quantities, dates, and locations. By harnessing GPT-4, we ensure that the templates are linguistically diverse and contextually rich, contributing to the overall quality and diversity of the dataset.
@@ 2.1.2 Simultaneous Q&A Generation and Verification
In a single integrated step, we generate specific problems and their corresponding solutions by substituting parameters into the GPT-4-generated meta-templates. Parameters are carefully selected to satisfy specific conditions, ensuring the solvability and validity of the problems. This simultaneous generation of questions and answers maintains consistency between the problem statements and their solutions.
To guarantee the correctness and reliability of the generated data, we employ a reject-sampling-based verification process. This involves executing the code-based solutions using a code executor and utilizing LLMs for solution verification. If a problem-solution pair fails verification—due to issues like runtime errors, incorrect results, or ill-formed language—it is discarded. Only verified and runnable templates are retained, ensuring the integrity of the dataset.
This iterative process of generation and verification continues until a sufficient number of high-quality problem-solution pairs are obtained. Integrating generation and verification into a single step streamlines the workflow and enhances the efficiency of the data generation process.
@@ 2.2 Process Flowchart
An illustrative overview of our TDG method is presented in Figure 1. The flowchart demonstrates the process starting from the LLM (e.g., GPT-4) generating meta-templates to the final dataset creation. After the meta-templates are generated, parameters are substituted to instantiate problems and solutions simultaneously. The generated Q&A pairs undergo verification using code execution and LLM-based checks. This loop continues, discarding any invalid pairs, until the dataset is populated with verified, high-quality data.
@@ 2.3 Code Implementation Example
An illustrative example of our TDG method is presented in Figure 2. The code snippet demonstrates how we generate problems involving sales over two consecutive months. The meta-template for this problem type was generated using GPT-4, capturing a realistic scenario that can be varied through parameter substitution. We include lists of random terms such as names, items, months, and locations to create diverse and realistic problem contexts. This randomness introduces variability and prevents the dataset from becoming repetitive, which helps in training models to generalize better.
@@ 2.4 Generated Problem and Solution Example
To illustrate the output of our TDG method, consider the following example generated using a GPT-4-produced template:
@@
Emily initially has 15 apples. She buys times more, so she buys apples. Now she has apples. She gives away 5 apples, so she is left with apples. Therefore, Emily has 55 apples now.
@@
This problem and solution pair were generated simultaneously and verified using code execution and LLM checks to ensure correctness.
@@ 2.5 Advantages of TDG
The TDG method offers several significant advantages that make it particularly effective for generating large-scale mathematical datasets:
Scalability: TDG enables the generation of an effectively infinite amount of data by varying parameters within GPT-4-generated templates. This scalability is crucial for training large language models that require vast amounts of data.
Quality Assurance: By integrating generation and verification into a single step with reject-sampling and utilizing code execution and LLM verification, we ensure that each problem-solution pair is correct and reliable. This precise supervision enhances the quality of the dataset and the performance of models trained on it.
Efficiency: The iterative process of generating and verifying Q&A pairs streamlines data creation, allowing for efficient accumulation of high-quality data.
Diversity: The use of GPT-4 to generate meta-templates introduces a wide variety of problem structures and linguistic styles, enhancing the diversity of the dataset. This diversity helps models generalize better to new and unseen problems.
Elevated Data Augmentation: By incorporating GPT-4 into the template generation process, we elevate data augmentation to a new level, enabling the synthesis of data that is both varied and high-quality.
@@ 3.3 Dataset Availability
TemplateGSM is publicly available and can be accessed at https://huggingface.co/datasets/math-ai/TemplateGSM. The code used for data generation is also provided at https://github.com/iiis-ai/TemplateMath. By sharing both the dataset and the generation code, we enable researchers and practitioners to reproduce our results, extend the dataset, and apply the TDG method to other domains or problem types.
@@ 4 Related Work
Our work builds upon these efforts by providing a significantly larger and more diverse dataset of mathematical problems with verified solutions, addressing the need for high-quality training data in mathematical reasoning.
@@
Base LLMs trained on vast corpora have demonstrated impressive language capabilities (Brown et al., 2020; Touvron et al., 2023). However, their performance on mathematical tasks is limited due to the scarcity of domain-specific training data. Recent studies have explored fine-tuning LLMs on mathematical datasets through continual pre-training (Lewkowycz et al., 2022; Azerbayev et al., 2023) or supervised fine-tuning (SFT) (Yu et al., 2023; Yue et al., 2023; Weng et al., 2023).
Continual pre-training involves further training on mathematical texts, enhancing models’ familiarity with mathematical language but not necessarily improving problem-solving skills. SFT approaches fine-tune models on curated question-answer pairs but are constrained by the availability of high-quality datasets.
Our TDG method enables the generation of extensive, high-quality problem-solution pairs, providing a valuable resource for both pre-training and fine-tuning LLMs, potentially leading to significant improvements in mathematical reasoning performance.
@@ 5 Conclusion
We have introduced Template-based Data Generation (TDG), a novel approach for generating large-scale, high-quality mathematical datasets through parameterized templates generated by GPT-4. Utilizing TDG, we created TemplateGSM, a dataset of over 7 million synthetically generated grade school math problems with verified solutions in both code and natural language formats.
Our extensive experiments demonstrate that TemplateGSM significantly enhances the mathematical reasoning capabilities of LLMs when used for pre-training and fine-tuning. The precise supervision offered by code execution and verification ensures the reliability of the data, fostering the development of models with improved understanding and problem-solving abilities.
@@
Limitations.
While TDG and TemplateGSM offer substantial benefits, there are limitations. One limitation is template bias, where models may become biased toward the structures present in the GPT-4-generated templates. Additionally, the generated problems are primarily at the grade school level, so extending TDG to higher-level mathematics requires careful template design, reflecting challenges in addressing complexity levels. Moreover, although GPT-4 introduces linguistic diversity, the style may still differ from human-authored educational materials, indicating a limitation in authenticity.
Future Work.
Future research directions include expanding template coverage by developing templates for more advanced mathematical topics, potentially leveraging even more sophisticated LLMs or collaborative human-AI template generation. Integrating augmentation techniques, such as using language models to rephrase and compose problems (Liu et al., 2024), can increase linguistic diversity further. Extending TDG to generate problems in multiple languages would create multilingual datasets. Additionally, exploring methods to mitigate template bias and enhance the authenticity of problem statements could improve the utility of the dataset. Conducting studies to assess the quality and educational value of the generated problems through human evaluation would provide valuable insights.

### v3

@@ Abstract
The rapid advancement of large language models (LLMs) such as GPT-3, PaLM, and Llama has significantly transformed natural language processing, showcasing remarkable capabilities in understanding and generating language. However, these models often struggle with tasks requiring complex reasoning, particularly in mathematical problem-solving, due in part to the scarcity of large-scale, high-quality, domain-specific datasets necessary for training sophisticated reasoning abilities. To address this limitation, we introduce Template-based Data Generation (TDG), a novel approach that leverages LLMs (GPT-4) to automatically generate parameterized meta-templates, which are then used to synthesize a vast array of high-quality problems and solutions. Leveraging TDG, we create TemplateMath Part I: TemplateGSM, a dataset comprising over 7 million synthetically generated grade school math problems—each accompanied by code-based and natural language solutions—with the potential to generate an effectively unlimited number more. This dataset alleviates the scarcity of large-scale mathematical datasets and serves as a valuable resource for pre-training, fine-tuning, and evaluating LLMs in mathematical reasoning. Our method not only enables the generation of virtually infinite data but also elevates data augmentation to a new level by using GPT-4 for meta-template generation, ensuring diverse and high-quality problem structures. The TemplateMath Part I: TemplateGSM dataset is publicly available at huggingface.co/datasets/math-ai/TemplateGSM222The code is available at github.com/iiis-ai/TemplateMath..
@@ 1 Introduction
We introduce TDG, a scalable method for generating an effectively infinite amount of high-quality, domain-specific data using parameterized templates generated by GPT-4. By leveraging GPT-4 to create meta-templates, we advance data augmentation to a new level, ensuring a diverse and rich set of problem structures for data synthesis.
@@ 2 Template-based Data Generation
Template-based Data Generation (TDG) is a systematic approach for producing large volumes of mathematical problems and corresponding solutions using parameterized templates. To advance data augmentation beyond traditional methods, we employ GPT-4 to generate meta-templates, broadly-defined structures capturing diverse problem types and linguistic styles. By varying parameters within these GPT-4-generated templates, TDG ensures scalability and maintains high-quality outputs, ultimately supporting the training and evaluation of large language models in mathematical reasoning tasks.
@@ 2.1 Methodology
The TDG process comprises several key components that collectively yield a high-quality mathematical dataset:
@@ 2.1.1 Generation of Meta-Templates with LLMs
We begin by leveraging large language models (LLMs), such as GPT-4, to generate meta-templates (Zhang et al., 2023). These meta-templates encode the underlying structures of various mathematical problem categories. GPT-4’s advanced language-generation capabilities allow us to produce a wide range of templates, each reflecting different mathematical concepts and problem types. Placeholders for variable components (e.g., names, items, quantities, dates, locations) ensure that the final dataset is both linguistically diverse and contextually rich, contributing to a comprehensive training corpus.
@@ 2.1.2 Simultaneous Q&A Generation and Verification
In a single integrated step, we generate specific problems and their corresponding solutions by substituting parameters into the GPT-4-produced meta-templates. Each parameter is carefully selected to satisfy certain problem requirements, ensuring both solvability and validity. The simultaneous generation of question and answer pairs preserves consistency between the problem statements and their solutions.
To guarantee correctness and reliability, we employ a rejection-sampling-based verification process. This entails running code-based solutions using a code executor and invoking LLMs for solution verification. Any problem-solution pair that fails—due to issues like runtime errors, incorrect results, or ill-formed language—is discarded. Only verified, executable templates are included in the final dataset, safeguarding data integrity.
This iterative generation-and-verification pipeline continues until an adequate number of high-quality problem-solution pairs is produced. By integrating both steps into one workflow, we streamline the data creation process and enhance overall efficiency.
@@ 2.2 Process Flowchart
An illustrative overview of our TDG method is shown in Figure 1. The flowchart begins with meta-template generation by an LLM (e.g., GPT-4) and concludes with the final dataset. After meta-templates are created, parameters are substituted to instantiate both problems and solutions. The resulting Q&A pairs then undergo verification via code execution and LLM-based checks. Any invalid pairs are discarded, and the process iterates until the dataset is populated with validated, high-quality data.
@@ 2.3 Code Implementation Example
An illustrative example of our TDG method is presented in Figure 2. The code snippet shows how we generate problems concerning sales over two consecutive months. The meta-template for this scenario, produced by GPT-4, can be varied through parameter substitution to yield many distinct problems. We maintain lists of random terms (e.g., names, items, months, locations) to ensure diversity and realism. This variation prevents repetition and helps models generalize better.
@@ 2.4 Generated Problem and Solution Example
To demonstrate TDG’s output, consider this example generated from a GPT-4-created template:
@@
Emily starts with 15 apples. She purchases times more, meaning apples. This brings her total to . After giving away 5 apples, she is left with . Thus, Emily has 55 apples remaining.
@@
This Q&A pair is generated and verified in a single step, with correctness confirmed via both code execution and LLM checks.
@@ 2.5 Advantages of TDG
TDG delivers several notable advantages for generating large-scale mathematical datasets:
Through parameter variation in GPT-4-generated templates, TDG can produce effectively infinite data, addressing the extensive training needs of large language models.
By integrating generation and verification into one process, and through code execution and LLM-based validation, every problem-solution pair is confirmed to be correct and consistent. This meticulous supervision boosts dataset quality and model performance.
GPT-4-generated meta-templates encompass a wide range of problem structures and linguistic variants, improving generalization to unseen problem types, and enabling the synthesis of large-scale, high-quality data.
@@ 3.3 Dataset Availability
TemplateGSM is publicly available and can be accessed at datasets/math-ai/TemplateGSM. The code used for data generation is also provided at github.com/iiis-ai/TemplateMath. By sharing both the dataset and the generation code, we enable researchers and practitioners to reproduce our results, extend the dataset, and apply the TDG method to other domains or problem types.
@@ 4 Related Work
Base LLMs trained on vast corpora have demonstrated impressive language capabilities (Brown et al., 2020; Touvron et al., 2023). However, their performance on mathematical tasks is limited due to the scarcity of domain-specific training data. Recent studies have explored fine-tuning LLMs on mathematical datasets through continual pre-training (Lewkowycz et al., 2022; Azerbayev et al., 2023) or supervised fine-tuning (SFT) (Yu et al., 2023; Yue et al., 2023; Weng et al., 2023). Continual pre-training involves further training on mathematical texts, enhancing models’ familiarity with mathematical language but not necessarily improving problem-solving skills. SFT approaches fine-tune models on curated question-answer pairs but are constrained by the availability of high-quality datasets.
@@ 5 Conclusion
We have introduced Template-based Data Generation (TDG), a novel approach for generating large-scale, high-quality mathematical datasets through parameterized templates generated by GPT-4. Utilizing TDG, we created TemplateGSM, a dataset of over 7 million synthetically generated grade school math problems with verified solutions in both code and natural language formats. The precise supervision offered by code execution and verification ensures the reliability of the data, fostering the development of models with improved understanding and problem-solving abilities.